# Dual-View Data Hallucination with Semantic Relation Guidance for Few-Shot Image Recognition

### v1

@@ I Introduction
The excellent performance of deep learning models [1, 2, 3, 4, 5] generally relies on a large amount of labelled data, which can be costly or infeasible to collect in diverse real-world applications. In contrast, human is able to learn new concepts from very few data samples. Motivated by this remarkable capacity of the human recognition system, few-shot learning (FSL) [6, 7, 8, 9, 10] has emerged as a popular research trend in recent years, which seeks to learn new concepts with minimal labeled data.
@@
Figure 1: t-SNE visualization of image samples from several classes and their semantic distance with the “ferret” class. Each class is denoted with a different color and class “ferret” is assumed to have one-shot sample (support set). Best viewed in color.
@@ II-1 Few-shot learning
Many existing FSL approaches follow the meta-learning paradigms, i.e., to learn a meta-learner from a lot of meta tasks that can quickly adapt to new few-shot tasks. In this work, we introduce the FSL works from a more concrete taxonomy and categorize them by four main perspectives: metric, optimization, model, and data. It is not a strict division since some works involve several perspectives. (i) Metric-based methods [11, 13, 21, 22, 23, 24, 6, 25, 26] mainly focus on learning an embedding model that maps the original data into a representation space suitable for few-shot recognition, where same-class samples are close while different-class samples are faraway under a specific metric. The Prototypical Network [22] is a classical method of this kind. It simply averages the few-shot embedded features of a novel class as the class prototype and recognizes a query sample by nearest-neighbor classifiers. The cosine metric is combined with deep neural networks for few-shot learning in the Matching Net [11], while the Mahalanobis distance is exploited in [27]. Zhang et al. [21] explore differentiable Earth Mover’s Distance as the metric function on top of deep networks. (ii) Optimization-based methods [15, 12, 28, 29, 30, 31, 32, 33, 34, 35, 36] aim to develop optimization algorithms that can train a good model quickly with few-shot samples. Finn et al. [12] propose a model-agnostic meta-learning algorithm to learn a good initialization for fast adaptation of deep networks in few-shot scenarios. Rusu et al. [31] propose to learn a low-dimensional latent embedding of model parameters and perform optimization-based meta-learning in this space. Neg-Cosine [30] uses negative margin loss to learn more robust backbone. In [36], Fei et al. utilize cross-episode optimization constraint between meta-learning episodes with the same set of classes. (iii) Model-based methods [37, 38, 14, 39, 40, 41] address the few-shot learning task by utilizing tailored models to generate the classifier parameters or designing sophisticated classification models. Lai et al. [14] learn a task-adaptive classifier-predictor to generate classifier weights by a meta-learner. Qiao et al. [37] propose to directly predict parameters from the activation statistics. In [38], a graph neural network denoising autoencoder model is developed to generate classification weights for few-shot learning. Liu et al. [39] propose to meta-learn an ensemble of epoch-wise empirical Bayes models for robust predictions in few-shot tasks. (iv) Data-based methods [10, 42, 43, 44, 45] overcome the lack of data by generating proper data for the novel classes directly. Hariharan and Girshick [42] learn in-class transformations from base classes and transfer them to novel classes by a multi-layer perceptron model. A delta-encoder is developed in [17] to extract transferable intra-class deformations, which are then applied to novel classes for sample synthesis. Yang et al. [18] uses statistics of base classes to calibrate the data distribution of novel classes and generate new data.
Zhang et al. [45] utilize a saliency network to obtain the foregrounds and backgrounds of image samples and hallucinate data from foreground-background combinations. Li et al. [10] propose an adversarial feature hallucination network that is built on a conditional Wasserstein Generative Adversarial Network (GAN) and takes the few labeled samples and random noise as input to hallucinate features. While most of these data-based methods solely exploit visual information, our work studies incorporating semantic relations to guide data generation from two views.
@@ III-C2 Semantic representation
Our proposed framework uses semantic information to guide generating hallucinated samples for training and involves semantic feature extraction from texts. In this work, we use DistilBERT [57] as the semantic feature extractor, which is a distilled version of BERT with lighter weight and faster computation speed. For each class $c$, we extract a semantic class vector $\mathbf{v}_{c}\in\mathbf{R}^{m}$ using the pre-trained DistilBERT model $\phi_{db}$, formulated as
$\mathbf{v}_{c}=\phi_{db}(S_{c}),$ (2)
where $S_{c}$ is the semantic label text for class $c$. To make better use of semantic information, we can use the class label to get its corresponding definition in WordNet [19] and combine the two as the semantic text.
@@ IV-A3 Implementation Details
We adopt WRN-28-10 [63] as the backbone, which is commonly used in few-shot learning tasks. For each dataset, the backbone is pre-trained on the base classes. The output feature dimension is 640. The classifier is learned using the scikit-learn toolbox [64]. In the IVDH module, we utilize Grad-CAM [58] to generate spatial attention maps for hallucinating the samples. We train the fusion network on the base classes for 100,000 iterations, using SGD optimizer with batch size of 5 and learning rate as 0.01. The parameter $\lambda$ is set as 0.3. For the PVDH module, the parameter $\tau$ for Turkey’s Ladder of Powers transformation is set as 0.5 by following previous works [59, 18]. We set $q$=2 to select similar base classes with $p$=10/32 for 1-shot/5-shot settings, and set $\alpha$=0.6 for prototype estimation. The number of resampled features is 200 per class.
@@ IV-D Ablation Study
Module analysis: Our proposed framework consists of two main modules, i.e., IVDH and PVDH. The baseline model of our method for classification is a WRN-28-10 backbone network with a classifier head, as described in Section IV-A. In ablation study, the baseline follows the same training procedure of our method except that our method generates hallucinated data for model training, since our method is dedicated to complementing the few-shot data appropriately. Table IV shows the evaluation results of different modules of our framework on the miniImageNet dataset. IVDH represents the model performance by adding our instance-view data hallucination module onto the baseline, and it is analogous for PVDH. As can be observed, the IVDH and PVDH modules improve the classification accuracy of the baseline by 1.21% and 5.46% on the 5-way 1-shot setting, respectively. As for the 5-way 5-shot setting, the performance gain is 4.11% and 6.19%. By integrating the IVDH and PVDH modules, our proposed method further boosts the model performance to 70.64% and 84.71% accuracy for 1-shot and 5-shot settings, yielding 5.93% and 6.37% performance gain from the baseline, respectively.
@@
Figure 4: Effect analysis of $\lambda$ and $\tau$.
Figure 5: Effect analysis of the parameters $\alpha$, $p$, and $q$ in PVDH.
@@
Figure 6: Effect analysis of the number of resampled data.

### v2

@@ I Introduction
The excellent performance of deep learning models [1, 2, 3, 4, 5] generally relies on a large amount of labelled data, which can be costly or infeasible to collect in diverse real-world applications. In contrast, human is able to learn new concepts from very few data samples. Motivated by this remarkable capacity of the human recognition system, few-shot learning (FSL) [6, 7, 8, 9, 10, 11] has emerged as a popular research trend in recent years, which seeks to learn new concepts with minimal labeled data.
@@
Figure 1: t-SNE visualization of image samples from several classes and their semantic distance with the “ferret” class. Each class is denoted with a different color and class “ferret” is assumed to have one-shot sample (support set). The “arctic fox” class is much closer to the “ferret” class in data distribution than the “stage” class, but it is not correctly reflected by the support sample of the “ferret” class in the visual space. In contrast, such relationships can be revealed from the semantic space. Best viewed in color.
@@ II-1 Few-shot learning
Many existing FSL approaches follow the meta-learning paradigms, i.e., to learn a meta-learner from a lot of meta tasks that can quickly adapt to new few-shot tasks. In this work, we introduce the FSL works from a more concrete taxonomy and categorize them by four main perspectives: metric, optimization, model, and data. It is not a strict division since some works involve several perspectives. (i) Metric-based methods [12, 14, 22, 23, 24, 25, 6, 26, 27] mainly focus on learning an embedding model that maps the original data into a representation space suitable for few-shot recognition, where same-class samples are close while different-class samples are faraway under a specific metric. The Prototypical Network [23] is a classical method of this kind. It simply averages the few-shot embedded features of a novel class as the class prototype and recognizes a query sample by nearest-neighbor classifiers. The cosine metric is combined with deep neural networks for few-shot learning in the Matching Net [12], while the Mahalanobis distance is exploited in [28]. Zhang et al. [22] explore differentiable Earth Mover’s Distance as the metric function on top of deep networks. (ii) Optimization-based methods [16, 13, 29, 30, 31, 32, 33, 34, 35, 36, 37] aim to develop optimization algorithms that can train a good model quickly with few-shot samples. Finn et al. [13] propose a model-agnostic meta-learning algorithm to learn a good initialization for fast adaptation of deep networks in few-shot scenarios. Rusu et al. [32] propose to learn a low-dimensional latent embedding of model parameters and perform optimization-based meta-learning in this space. Neg-Cosine [31] uses negative margin loss to learn more robust backbone. In [37], Fei et al. utilize cross-episode optimization constraint between meta-learning episodes with the same set of classes. (iii) Model-based methods [38, 39, 15, 40, 41, 42] address the few-shot learning task by utilizing tailored models to generate the classifier parameters or designing sophisticated classification models. Lai et al. [15] learn a task-adaptive classifier-predictor to generate classifier weights by a meta-learner. Qiao et al. [38] propose to directly predict parameters from the activation statistics. In [39], a graph neural network denoising autoencoder model is developed to generate classification weights for few-shot learning. Liu et al. [40] propose to meta-learn an ensemble of epoch-wise empirical Bayes models for robust predictions in few-shot tasks. (iv) Data-based methods [11, 43, 44, 45, 46] overcome the lack of data by generating proper data for the novel classes directly. Hariharan and Girshick [43] learn in-class transformations from base classes and transfer them to novel classes by a multi-layer perceptron model. A delta-encoder is developed in [18] to extract transferable intra-class deformations, which are then applied to novel classes for sample synthesis. Yang et al. [19] uses statistics of base classes to calibrate the data distribution of novel classes and generate new data. Zhang et al. [46] utilize a saliency network to obtain the foregrounds and backgrounds of image samples and hallucinate data from foreground-background combinations. Li et al. [11] propose an adversarial feature hallucination network that is built on a conditional Wasserstein Generative Adversarial Network (GAN) and takes the few labeled samples and random noise as input to hallucinate features. While most of these data-based methods solely exploit visual information, our work studies incorporating semantic relations to guide data generation from two views.
@@ III-C2 Semantic representation
Our proposed framework uses semantic information to guide generating hallucinated samples for training and involves semantic feature extraction from texts. Let $\phi_{sr}$ denote the semantic feature extractor. For each class $c$, we extract a semantic class vector $\mathbf{v}_{c}\in\mathbf{R}^{m}$ using the pre-trained semantic feature extraction model $\phi_{sr}$, formulated as
$\mathbf{v}_{c}=\phi_{sr}(S_{c}),$ (2)
where $S_{c}$ is the semantic label text for class $c$.
@@ IV-A3 Implementation Details
We adopt WRN-28-10 [63] as the backbone (visual feature extractor), which is commonly used in few-shot learning tasks. The backbone ends with ReLU activation. For each dataset, the backbone is pre-trained on the base classes. The output feature dimension is 640. The classifier is learned using the scikit-learn toolbox [64]. For the semantic feature extractor $\phi_{sr}$, we employ DistilBERT [65], which is a distilled version of BERT [66] with a reduced weight and faster computation speed while retaining comparable performance. To make better use of semantic information, we use the class label to get its definition in WordNet [20] and combine the two as the semantic text $S_{c}$ if the corresponding definition exists. In the IVDH module, we utilize Grad-CAM [58] to generate spatial attention maps for hallucinating the samples. We train the fusion network on the base classes for 100,000 iterations, using SGD optimizer with batch size of 5 and learning rate as 0.01. The parameter $\lambda$ is set as 0.3. For the PVDH module, the parameter $\tau$ for Turkey’s Ladder of Powers transformation is set as 0.5 by following previous works [59, 19]. We set $q$=2 to select similar base classes with $p$=10/32 for 1-shot/5-shot settings, and set $\alpha$=0.6 for prototype estimation. The number of resampled features is 200 per class.
@@ IV-D Ablation Study
Module analysis: Our proposed framework consists of two main modules, i.e., IVDH and PVDH. The baseline model of our method for classification is a WRN-28-10 backbone network with a classifier head, as described in Section IV-A. In ablation study, the baseline follows the same training procedure of our method except that our method generates hallucinated data for model training, since our method is dedicated to complementing the few-shot data appropriately. Table IV shows the evaluation results of different modules of our framework on the miniImageNet dataset. IVDH represents the model performance by adding our instance-view data hallucination module onto the baseline, and it is analogous for PVDH. As can be observed, the IVDH and PVDH modules improve the classification accuracy of the baseline by 1.21% and 5.46% on the 5-way 1-shot setting, respectively. As for the 5-way 5-shot setting, the performance gain is 4.11% and 6.19%. By integrating the IVDH and PVDH modules, our proposed method (Ours-full) further boosts the model performance to 70.64% and 84.71% accuracy for 1-shot and 5-shot settings, yielding 5.93% and 6.37% performance gain from the baseline, respectively.
We also conduct an ablation experiment on the effect of the semantic text $S_{c}$. While Ours-full uses the combination of the label and its definition in WordNet, let Ours-l denote a variant of our method that only uses the label for $S_{c}$. As observed from Table IV, the two perform comparably, with Ours-full showing slightly better performance by using more semantic information.
@@
Figure 4: Effect analysis of $\lambda$ and $\tau$ on miniImageNet.
Figure 5: Effect analysis of the parameters $\alpha$, $p$, and $q$ in PVDH on miniImageNet.
@@
Figure 6: Effect analysis of the number of resampled data on miniImageNet.
@@ IV-E Visualization Analysis
IV-F Discussion and Potential Extensions
In the previous subsections, extensive experimental comparisons and analyses have shown the effectiveness of our method. However, certain limitations may arise in some cases, as we observed with the Omniglot dataset (illustrated in Figure 3), where images and labels offer limited semantic information, leading to less satisfactory performance of our method. In future work, we will investigate potential solutions to overcome these limitations and improve the generalizability of our method. Additionally, there are several other aspects of our method that merit further exploration. Firstly, in the prototype-view data hallucination module, we compute the semantic and visual similarity in their respective domains and combine them via ranking to model the semantic-visual relationship. It is worth exploring better semantic-visual relation modeling schemes, such as directly bridging the two domains as in CLIP [78]. Secondly, we plan to investigate the scalability of our proposed method beyond images and explore its potential for few-shot video classification [79]. Actually, our method can be directly applied to this task by replacing image features with video features. However, the complex spatio-temporal context of videos may pose challenges, and developing proper mechanisms to better exploit such context may be needed for effective hallucination. While Gao et al. [80] introduce an innovative method to learn prototypes from a group of semantic word vectors for zero-shot video classification, we plan to incorporate visual information to build more effective prototypes of video categories for data hallucination in few-shot scenarios, thereby better exploiting the advantages of our method.